A recent study shows that the annual cost of maintaining a building in Accra averages
200 Cedis with a variance of 260 Cedis. If a tax of 30% is introduced on all items
associated with the maintenance of building i.e. everything is made 30% more expensive.
Calculate the standard deviation of the annual cost of maintaining a building in Accra
1
Expert's answer
2020-05-13T18:56:39-0400

"\text {Let Z represents the annual cost of maintaining}\\
\text { a building in Accra, then, E(Z)=200, V(Z)=260,}\\
\text {Let X represents the annual cost of maintaining}\\
\text { a building in Accra after adding a tax of 30\%,}\\
X=1.30Z\\
V(X)=V(1.30Z)=1.30^{2}V(Z)\\
=1.30^{2}(260)=439.4,\\
\therefore \sigma_{X}=\sqrt{V(X)}=\sqrt{439.4}=20.962"